OpenStudy (lukecrayonz):

Find an equation for the hyperbola Vertices: (2, 3), (2, -3); Foci: (2, 5), (2, -5)

OpenStudy (anonymous):

This is gonna be a vertical hyperbola since we have change in the y coordinates, meaing the y comes first in the eqaution. Also this is a shifted hyperbola (duh)

OpenStudy (anonymous):

hold on man, i just totally blanked out

OpenStudy (lukecrayonz):

Heres an example for you http://screensnapr.com/v/eRu1wi.png

OpenStudy (mertsj):

c=2, a=3 a^2+b^2=c^2

OpenStudy (lukecrayonz):

b^2=5

OpenStudy (lukecrayonz):

Then finding the center, what coordinates do we use?

OpenStudy (mertsj):

\[\frac{(y-0)^2}{9}-\frac{(x-2)^2}{5}=1\]